### v1

# WRDScore: New Metric for Evaluation of Natural Language Generation Models

Abstract.
The problem of natural language generation, and, more specifically, method name prediction, faces significant difficulties when proposed models need to be evaluated on test data. Such a metric would need to consider the versatility with which a single method can be named, with respect to both semantics and syntax. Measuring the direct overlap between the predicted and reference (true) sequences will not be able to capture these subtleties. Other existing embedding based metrics either do not measure precision and recall or impose strict unrealistic assumptions on both sequences. To address these issues, we propose a new metric that, on the one hand, is very simple and lightweight, and, on the other hand, is able to calculate precision and recall without resorting to any assumptions while obtaining good performance with respect to the human judgement.
software engineering, recommender systems, method name prediction
1. Introduction
The current rapid rise of deep-learning-based natural language generation models acutely necessitates the existence of smart metrics to evaluate these emerging models. Nowadays, the available metrics are unable to capture more subtle differences in semantics and syntax between generated and reference texts. That is, such simple and widely accepted metrics as ROUGE (Lin, 2004) can measure only the direct overlap between the tokens in reference and generated sequences. This naive approach does not consider possible paraphrasing, replacement of some tokens with synonyms, abbreviations, the order of tokens, etc. For instance, ROUGE-1 metric would not be able to detect the obvious similarity between words “copy” and “clone”, or “size” and “get count”.
In this paper, our goal is to create a novel metric for evaluation of the output of natural language generation (NLG) models, which is better tailored to the prediction problem of Java method names. Other embedding-based NLG metrics, such as BERTScore (Zhang et al., 2020), make unnecessary implicit assumptions. For instance, in the framework of BERTScore, one token from the reference sequence can be matched only to a single token in the predicted sequence, and vice versa. In this paper, we will show that this assumption can not only be safely relaxed, but leads to the considerable improvement even without using the advanced deep contextualized embeddings, such as BERT (Devlin et al., 2019).
To overcome the above-mentioned impediments, we propose our new NLG metric — Word Rotator’s Distance Score (WRDScore). One of the distinguishing features of WRDScore is that its output value is normalized to lie between 0 and 1. Such values provide more meaning than non-normalized ones. A normalized score is more interpretable and conveys a better sense of how close an NLG model performs to the best possible case. Moreover, WRDScore can be powered by any kind of word embeddings, including contextualized (ELMo, BERT) and non-contextualized ones (Word2Vec, GloVe). Leveraging the optimal transport formulation in WRDScore leads to a soft assignment between the compared sentences. Thus, this technique introduces context into measurements while somewhat relaxing the dependence of the resulting score on the context encoded in the used word embeddings. This natural property allows WRDScore to outperform other advanced NLG metrics, such as BERTScore, even without resorting to contextualized word embeddings. In this sense, WRDScore can be seen as a generalization and relaxation of BERTScore.
2. Related work
2.1. ROUGE
ROUGE metric (Lin, 2004), or Recall-Oriented Understudy for Gisting Evaluation, is the simplest and most widely used approach to compare two text sentences. It relies on the direct overlap between the two token sets of the input sentences. The most popular variation of ROUGE is the ROUGE-N metric, with N equal to 1 or 2. More precisely, ROUGE-N measures the number of matching n-grams the model-generated text and a human-produced reference sentence. ROUGE-N is a recall-oriented measure defined as follows:
$ROUGE-N=\frac{\sum_{S\in\{ReferenceSummaries\}}\sum_{gram_{n}\in S}Count_{match}(gram_{n})}{\sum_{S\in\{ReferenceSummaries\}}\sum_{gram_{n}\in S}Count(gram_{n})}$
Here $n$ stands for the length of the n-gram, $gram_{n}$, and $Count_{match}(gram_{n})$ is the maximum number of n-grams co-occurring in a candidate summary and a set of reference summaries. A closely related measure, BLEU, used in automatic evaluation of machine translation, is a precision-based measure. BLEU measures how well a candidate translation matches a set of reference translations by counting the percentage of n-grams in the candidate translation overlapping with the references.
2.2. BERTScore
BERTScore (Zhang et al., 2020) uses BERT (Devlin et al., 2019) contextualized embeddings to compute pairwise cosine similarity between each hypothesis token $i$ with each reference token $j$. Then, the precision and recall are calculated by maximum-based hard matching of tokens in the hypothesis sentence to the tokens of the reference sentence as follows:
$BERTScore=F_{BERT}=2\frac{P_{BERT}\cdot R_{BERT}}{P_{BERT}+R_{BERT}}$
$R_{BERT}=\frac{1}{|r|}\sum_{i\in r}\max_{j\in p}\vec{i}^{T}\vec{j}$
$P_{BERT}=\frac{1}{|p|}\sum_{j\in p}\max_{i\in r}\vec{i}^{T}\vec{j}$
2.3. MoverScore
MoverScore (Zhao et al., 2019) was inspired by using the Word Mover’s Distance (Kusner et al., 2015) (WMD) and takes the value equal to the cost of the optimal transportation plan between the reference and predicted sentences when they are viewed as probability distributions. MoverScore is able to use Word2Vec, BERT, or ELMo to define word embeddings. Also, MoverScore assigns weights for the defined word vectors via TF-IDF. Word embeddings can be n-grams.
It was observed that MoverScore (Zhao et al., 2019) only marginally outperforms BERTScore when measured on a variety of tasks, such as machine translation, image captioning, and dialog response generation. MoverScore outperforms both asymmetric HMD (“Hard Mover’s Distance”) factors (Zhao et al., 2019). However, if they are combined via harmonic mean, BERTScore is on par with MoverScore. We conjecture that BERT softens hard alignments of BERTScore as contextualized embeddings encode information about the whole sentence into each word vector.
BERTScore is a precision-recall-based metric. Thus, it is naturally normalized and provides a better interpretation of an NLG model’s performance. This interpretation is useful when the model’s predictions either lack precision (i. e. do not guess true tokens precisely) or recall (do not cover true sentences well enough), or the evaluation dataset is highly imbalanced. The vast majority of embedding-based NLG metrics proposed in the literature neither follow the precision-recall framework, nor output normalized scores. In contrast, we propose a natural mean-based approach to measure precision and recall from the optimal transport flow values between the reference and predicted sentences. To the best of our knowledge, no other existing embedding-based metric provides such a property. Also, contrary to MoverScore, the word vectors used in our algorithm are pre-normalized and weighted by TF-IDF scores. This enables to consider the pairwise cosine similarity between word vectors instead of Euclidean distance used in MoverScore.
3. Proposed Algorithm
We will describe a step-by-step process of building the proposed WRDScore metric. First, for a given vocabulary of tokens $V=\{v_{1},\ldots,v_{n}\}$, we train word embeddings by means of a very simple co-occurrence method. A square cooccurrence matrix $C=\{c_{ij}\}\in\mathbb{R}^{n\times n}$ is filled, where each element $c_{ij}$ indicates how many times token $i$ has occurred with token $j$ in one method name of the training corpus. Since this matrix is very large and sparse to be store in memory, we approximate it with a lower-dimensional matrix $E\in\mathbb{R}^{n\times e}$ via the singular value decomposition (SVD). Here $e\in\mathbb{N}$ is the embedding dimension, which was chosen to be $50$ in our experiments. Matrix $C$ is decomposed as follows:
$C\stackrel{{\scriptstyle SVD}}{{=}}U\cdot\Sigma\cdot V=E\cdot 2V$
where $E=\frac{1}{2}\cdot U\cdot\Sigma$ is the embedding matrix: each of its rows is the embedding vector of the correspoding token from the vocabulary.
This simple cooccurrence-based approach for building token embeddings is not the only choice. Some more advanced methods can be also be employed, such as Word2Vec (!!!cite!!!) or BERT. During our experiments, we observed that training Word2Vec over the training corpus with the minimal context window size, which is 2, results in embeddings that are able to capture synonym relations between tokens most precisely. This leads to a better overall performance of WRDScore.
Suppose a reference and predicted sequences are given by $r=\{t_{1},\ldots,t_{l}\}$ and $p=\{\hat{t}_{1},\ldots,\hat{t}_{m}\}$, respectively. Then, the usual notions of precision and recall can be calculated as follows:
$TP=|r\cap p|$
$FP=|p\setminus r|$
$FN=|r\setminus p|$
$PR=\frac{TP}{TP+FP}\ \ \text{precision}$
$RC=\frac{TP}{TP+FN}\ \ \text{recall}$
If we associate each token with its one-hot embedding vector, i.e. $t_{i}\mapsto(0,\ldots,0,1,0,\ldots,0)$, where $1$ is placed at the $i^{th}$ position, then the above precision and recall formulas are just special cases of a more general formula:
(1) $PR=\frac{1}{|p|}\sum_{i\in p}\max_{j\in r}\vec{i}^{T}\cdot\vec{j}$
(2) $RC=\frac{1}{|r|}\sum_{j\in r}\max_{i\in p}\vec{i}^{T}\cdot\vec{j}$
Also, we can use these formulas by substituting one-hot embeddings with more advanced ones, like the embeddings obtained from the cooccurrence matrix. However, in these formulas, we assume that each token vector in the reference sequence is associated with a single token vector in the prediction sequence, which clearly might not hold in general: a single token from the reference can correlate with multiple words in the prediction, and vice versa.
To overcome the above limitation, we propose to use Word Rotator’s Distance (Yokoi et al., 2020). Now, suppose that the reference and predicted sequences are equipped with some set of weights $W=\{w_{1},\ldots,w_{l}\}$ and $\hat{W}=\{\hat{w}_{1},\ldots,\hat{w}_{m}\}$, respectively. The embedding norms of the corresponding tokens are usually taken as weights. This choice is justified since the norm of a token embedding is usually proportional to the token’s importance in the training corpus. These weights can also be augmented with the TF-IDF coefficients. Next, let $\{F_{ij}\}$, $i=1,\ldots,l$, $j=1,\ldots,m$, be the flow matrix which was the result of the solution of the following optimization problem:
(3) $\min_{F\in\mathbb{R}^{l\times m}}\sum_{i,j}F_{ij}\cdot d_{ij}$
subject to
$F\cdot\mathbf{1}=W$
$\mathbf{1}^{T}\cdot F=\hat{W}$
In the above equations, the distance is $d_{ij}=1-\cos(t_{i},\hat{t}_{j})$, where $\cos(\cdot,\cdot)$ denotes the cosine similarity between two vectors. Then, the Word Rotator’s Distance between two sequences is:
$WRD((r,W),(p,\hat{W}))=\sum_{i,j}F_{ij}\cdot d_{ij}$
Our goal is to devise a way to calculate the new generalized precision, recall, and F1 scores more accurately than in formulas 1 and 2 by taking into account the flow matrix $F_{ij}$, which expresses how closely token $i$ is related to token $j$. In this setting, one token can be associated not only with a single, but with multiple other tokens at once. We can leverage this information and define the precision and recall as follows:
(4) $PR=\frac{1}{|p|}\sum_{i\in p}\frac{\sum_{j\in r}F_{ij}\cdot\cos(t_{i},\hat{t}_{j})}{\sum_{j\in r}F_{ij}}$
(5) $RC=\frac{1}{|r|}\sum_{i\in r}\frac{\sum_{j\in p}F_{ij}\cdot\cos(\hat{t}_{j},t_{i})}{\sum_{j\in p}F_{ij}}$
The use of cosine distance in the optimization problem 3 and cosine similarity in 4 and 5 ensure that the similarity scores between pairs of tokens are naturally bound to be between 0 and 1. The use of Euclidean distance would not let us define the new precision and recall formulas without some artificial bounding procedure, whose optimality would need to be investigated. Thus defined metric is named Word Rotator’s Distance Score (WRDScore).
4. Experiments
To demonstrate the efficiency of the proposed metric, we conducted some experiments on the following large dataset of Java code. By training the embeddings via the simple cooccurrence model with the largest context window size on Java method names from this dataset, we evaluate the output of an NLG model, called UniXcoder (Guo et al., 2022), and obtain the following results. Each line below represents the reference and predicted sequences as well as the new recall, precision, and F1 scores, for which the old ROUGE-1 score was zero:
  1. (1)
    $r$ = “update image”, $p$ = “set artwork entry”, $pr$ = 0.5787, $rc$ = 0.4652, $f1$ = 0.5158
  2. (2)
    $r$ = “advance”, $p$ = “next char”, $pr$ = 0.7152, $rc$ = 0.7112, $f1$ = 0.7132
  3. (3)
    $r$ = “size”, $p$ = “get count”, $pr$ = 0.7669, $rc$ = 0.7597, $f1$ = 0.7632
  4. (4)
    $r$ = “receive message”, $p$ = “log”, $pr$ = 0.7939, $rc$ = 0.7720, $f1$ = 0.7828
  5. (5)
    $r$ = “advance”, $p$ = “next char”, $pr$ = 0.7152, $rc$ = 0.7112, $f1$ = 0.7132
  6. (6)
    $r$ = “perform validation”, $p$ = “validate url”, $pr$ = 0.7849, $rc$ = 0.7877, $f1$ = 0.7863
  7. (7)
    $r$ = “do until result”, $p$ = “execute”, $pr$ = 0.5904, $rc$ = 0.5900, $f1$ = 0.5902
  8. (8)
    $r$ = “detect capital use”, $p$ = “is all capitalized”, $pr$ = 0.5672, $rc$ = 0.5225, $f1$ = 0.5439
  9. (9)
    $r$ = “longest consecutive”, $p$ = “max path”, $pr$ = 0.6778, $rc$ = 0.6627, $f1$ = 0.6702
  10. (10)
    $r$ = “schedule course”, $p$ = “get number of days”, $pr$ = 0.6409, $rc$ = 0.6351, $f1$ = 0.6380
  11. (11)
    $r$ = “process username”, $p$ = “verify”, $pr$ = 0.6930, $rc$ = 0.6878, $f1$ = 0.6904
We also conduct a human evaluation of the proposed metric by training token embeddings using a more advanced Word2Vec model with the size of context window equal to 2. We considered 3 sets of method name pairs consisting of 50 samples each, where the samples were obtained from the results generated by the UniXcoder model trained on the above-mentioned dataset of Java methods. The samples in each set were chosen as follows:
  1. (1)
    10 first samples were drawn uniformly at random from the entire dataset;
  2. (2)
    10 subsequent samples were drawn uniformly at random from the set of the method name pairs, for which the WRDScore metric is within the range from 00 to $0.2$;
  3. (3)
    10 next samples were drawn uniformly at random from the set of method name pairs, for which the WRDScore metric is within the range from $0.2$ to $0.5$;
  4. (4)
    10 subsequent samples were drawn uniformly at random from the set of method name pairs, for which the WRDScore metric is within the range from $0.5$ to $0.8$;
  5. (5)
    10 last samples were drawn uniformly at random from the set of method name pairs, for which the WRDScore metric is within the range from $0.8$ to $1.0$ and the ROUGE-1 metric is less than $0.5$ (imposing the latter range was needed to prevent mostly sampling identical name pairs).
We asked 9 experienced Java programmers to assign a semantic similarity score to each of the method name pairs in the range from 00, which means totally different, to $1.0$, which means totally similar, that is expessing method names of the same semantic meaning. Then, to account for the bias in perception of semantic similarity between different responders, we average the responses of 3 different people for each set. Then, considering these scores as the ground truth, we compute how far the corresponding ROUGE-1 and WRDScore metric values for each name pair deviate from the ground truth using the mean squared error (MSE) and mean absolute error (MAE). The following results have been obtained:
ROUGE-1 WRDScore p-value
MSE 0.0817 0.0678 $1.42\cdot 10^{-18}$
MAE 0.2213 0.1891 $5.07\cdot 10^{-37}$
Table 1. The observed mean deviation of the ground truth (human scores) from the scores obtained by the ROUGE-1 and WRDScore metrics
The reported values have been obtained by averaging the results of $100$ bootstraps and the p-values were calculated using T-tests for the means of two independent samples of scores. Thus, WRDScore was by $17.01$% better over ROUGE-1 with respect to MSE, and by $14.55$% better with respect to MAE.
5. Conclusion
In this paper, we have proposed a new metric for evaluation of natural language generation models. The precise details of the contruction of a new metric have been described. The novelty of our approach lies in releasing the assumption that in computing precision and recall of a generator one token from a true sequence can be matched to only one token in the predicted sequence, and vice versa. The new metric has been tested under two different token embedding schemes, and the obtained results clealy shows the superiority of the novel WRDScore metric over the competitive ones.
Here are some possible directions for future research:
  1. (1)
    Study how sensitive the proposed metric is to small changes in its inputs. It should vary continuously with respect to its input;
  2. (2)
    Investigate how more advanced token embeddings could enhance the proposed metrics;
  3. (3)
    Enhance the model in a way that would allow to take into account the order of tokens in reference and predicted sequences;
  4. (4)
    Study how the proposed metric can be effectively incorporated into the loss function used for training different NLG models.